# Do Large Language Models Have an English “Accent”? Evaluating and Improving the Naturalness of Multilingual LLMs

### v1

@@ 1 Introduction
We could make an analogy between these multilingual LLMs and native English speakers who are trying to acquire a new language (Groves et al., 2018; DeVore and Kyle, 2023). Their language notions are built in an English-centric system, and they inevitably bring traces of English habits into other languages when transferring their notions (Papadimitriou et al., 2023; Wendler et al., 2024). Moreover, because of the lack of data in non-English languages, multilingual LLMs are often exposed – either during pre-training or post-training (Yang et al., 2024; Abdin et al., 2024) – to texts translated from English. Both human and machine translated language are known to suffer from translationese artifacts, which set them apart from native content (Bizzoni et al., 2020; Luo et al., 2024). LLMs trained on such data are susceptible to suffer from the same unnaturalness problems as the second-language learners in our analogy.
@@ 2 Related Work
Multilingual analysis of LLMs has recently shown that models trained on unbalanced, English-heavy corpora often rely on English as an internal pivot language. Wendler et al. (2024) demonstrate that the concept space of LLaMA-2 is more closely aligned with English than with other input languages. Similarly, Papadimitriou et al. (2023) show that multilingual BERT exhibits a bias toward English-like grammatical structures. Despite these findings, there has been no systematic study on how this English-centric tendency affects the linguistic naturalness of multilingual LLM outputs, particularly in open-ended downstream tasks.
@@ 3 Evaluation Metrics for Naturalness
In this section, we present a new set of evaluation metrics aimed at measuring the naturalness of multilingual text generations at the corpus level. While our approach also requires a reference set of natively written texts, it differs from widely used reference-based metrics such as BLEU (Papineni et al., 2002), ROUGE (Lin, 2004), and BERTScore (Zhang* et al., 2020). Indeed, instead of comparing reference and system outputs at the sample level, our new metrics allow us to investigate naturalness at the corpus (distribution) level.
Sample-level reference metrics often struggle to account for human label variability and uncertainty (Plank, 2022; Giulianelli et al., 2023), particularly in open-ended tasks with multiple valid generations. Also, while a single text sample with certain choices of vocabulary or grammatical structure might seem natural, the repeated occurrence of these features across a large scale of generations would raise a red flag (Guo et al., 2023). That is, detecting unnaturalness in individual samples may prove challenging, but when analyzed at the corpus level, statistical features become strong indicators (Liang et al., 2024a, b). Our metrics operate by comparing patterns of language use across large text collections, providing a broader, distribution-based perspective.
Having highlighted the advantage of a corpus-level perspective, we propose a new definition for language model naturalness. Past studies have defined the naturalness of a piece of text by asking “could it have been produced by a native speaker?” (Novikova et al., 2016; Groves et al., 2018; Liu et al., 2021). We adapt this definition to the corpus level and define the naturalness of a language model by asking “could the set of texts generated by this language model have been produced by a group of native speakers when responding to the same prompts?”
Our metrics are inspired by other divergence-based measures such as MAUVE (Pillutla et al., 2021; Pimentel et al., 2023), which quantify the information-theoretic divergence between the probability distributions of a language generator and a true natural language distribution. However, our approach differs by not relying on another language model to embed the generated texts. This reduces the risk of introducing intrinsic biases from the chosen embedding model, a crucial consideration in multilingual settings, where such the chosen embedding models are often English-dominated themselves. Our method is also more transparent and interpretable, as it clearly distinguish between two key aspects of linguistic naturalness: syntactic and lexical naturalness. However, our metrics focus exclusively on the linguistic form of the text and do not address semantic aspects. For analyzing semantics, the use of external embedding models may be inevitable.
@@ 3.1 Lexical Naturalness
We propose to measure the lexical naturalness of an LLM by comparing the vocabulary distribution of its generated text with the that of human-written text. More specifically, we put forward a lexical naturalness metric based on computing the Jensen-Shannon Divergence (JSD) between the lexical distributions of LLM-generated and human-written text from the same prompts. The JSD provides a symmetric and bounded measure of difference between two distributions without them necessarily sharing the same support, i.e., the tokens in the vocabulary of the LLM in our case. [^2] We process vocabularies at the word level instead of subword/token level to represent more meaningful lexical units. Given the vocabulary distributions $P$ and $Q$ corresponding to human and model outputs, respectively, the JSD is calculated as follows:
@@ 3.2 Syntactic Naturalness
To measure the syntactic naturalness of LLMs generations we leverage the Universal Dependencies (UD) grammar framework (Nivre et al., 2020). UD provides well-defined, theoretically-grounded linguistic structures across different languages, making it suitable for cross-lingual comparisons. Our proposed metric is based on representing each sentence as a dependency tree, where nodes correspond to words and edges specify the dependency relations between them. Additionally, each word is annotated with its corresponding Part-of-Speech (POS) tag. At a high-level, our metric computes the structural similarity of all pairs of sentences in a corpus, clustering texts whose syntax is similar and allowing us to determine if there is a distributional difference between two groups (LLM-generated text and human-written text, in our case).
@@
Given a set of human-generated sentences $\{s_{i}^{\text{h}}\}_{i=1}^{N_{\text{h}}}$ and model-generated sentences $\{s_{j}^{\text{m}}\}_{j=1}^{N_{\text{m}}}$, we compute the WL kernel between all pairs of dependency trees from these sets. The resulting kernel matrix $\mathbf{K}\in\mathbb{R}^{N_{\text{h}}\times N_{\text{m}}}$ has elements $K_{ij}=K_{\text{WL}}(T_{i}^{\text{h}},T_{j}^{\text{m}})$, where $T_{i}^{\text{h}}$ and $T_{j}^{\text{m}}$ represent the dependency trees corresponding to sentences $s_{i}^{\text{h}}$ and $s_{j}^{\text{m}}$, respectively. This kernel matrix $\mathbf{K}$ captures structural similarities between the dependency trees of human-written and LLM-generated sentences.
@@
Our proposed syntactic naturalness metric quantifies the syntactic divergence between human and model-generated sentences by examining the distribution of their dependency tree structures. This approach considers both the dependency relationships and the hierarchical arrangement of substructures at multiple levels.
@@ 4 Cross-lingual Analysis of Naturalness
After introducing our new metrics for evaluating lexical and syntactic naturalness, we proceed with a cross-lingual analysis of the naturalness of multilingual LLMs. This process involves curating a new dataset, selecting the models, and analyzing the benchmark results.
@@ 4.1 Dataset and Evaluation
For our analysis, we require a corpus that is topically aligned across languages while preserving ground-truth naturalness. This means that the texts in each language must be natively written, not translated – whether by humans or models. As a result, the parallel corpora typically used in machine translation research are not suitable. Instead, we construct a new dataset that satisfies our criteria starting from Wikipedia [^3] Wikipedia is available under the CC BY-SA 4.0 license., which offers a wealth of texts that are frequently edited by native speakers and that we can topically-align across languages. Importantly, to minimize cultural bias in our dataset, we select the most-viewed Wikipedia articles across English, Chinese and French. [^4] We use the Wikipedia API to retrieve the top-20 most-viewed articles in each language daily from June 2022 to June 2025. We then take the per-language union of the articles and retain their the cross-language intersection. Our data curation process results in 3,722 Wikipedia pages, each accompanied by their text in the three target languages. We discuss the preprocessing of this dataset in Appendix A.
For each entry in the dataset, we instruct the models to complete a straightforward task: generating a description for the given entry in each of the three languages. We selected this task to avoid overly constraining the content of the language model outputs, allowing them to generate in a more natural and organic manner. We stress that our focus is on the overall distribution of vocabulary and grammatical structures rather than on individual outputs, the semantic content and/or factual accuracy. The prompts and generation settings used in this task are provided in Appendix B.
@@ 4.3 Results and Analysis
In terms of lexical divergence, English outputs – especially from Mistral-v0.3 – approach human reference values, while larger gaps persist in Chinese and French, with Chinese exhibiting the most pronounced difference. This suggests that models are lexically less natural in languages other than English.
@@
Previous benchmark results are generated using temperature sampling with $t=0.6$, which is found to provide the optimal balance between creativity and consistency for multilingual generation tasks (Dubey et al., 2024). As an ablation study, we vary the decoding temperature between 0.3 and 0.9 to examine its impact on the naturalness of model outputs. Our results in Figure 4 show that for syntax, natural models become more natural as the decoding temperature increases, while unnatural models exhibit increased unnaturalness. For vocabulary, all models show improved naturalness as the decoding temperature rises. It is also worth noting that the ranking of the models generally remains unchanged across different decoding settings.
@@ 5 Improving Naturalness through Preference Tuning
We now propose an approach to improve the naturalness of language model outputs. Among the various stages of LLM development, preference-based learning is the most effective for refining stylistic features Ivison et al. (2024). Based on this, we focus on aligning models for better naturalness during the preference tuning stage. Among the various preference optimization techniques, we propose to use DPO due to its simplicity and efficiency.
@@ 5.1 Preference Dataset Construction
We construct the preference datasets starting with SFT (Supervised Fine-Tuning) datasets originally written in the target language. However, we find that non-translated, open-source SFT datasets for non-English languages are almost non-existent. Even for Chinese – despite being the second highest-resource language after English – only a limited number is available. Consequently, we conduct our DPO experiments on Chinese, using two datasets that focus on essay generation and open-domain QA tasks. These tasks are well-suited for stylistic alignment due to their open-ended, creative nature. For essay generation, we use the “composition” split from the Firefly dataset (Yang, 2023), and for open-domain QA, we use the OpenLabel-Chinese Conversations Dataset  [^5] Released under the Apache 2.0 license. (BAAI, 2023).
The original instructions from the SFT datasets are preserved, with the initial response serving as the preferred response. To generate the unnatural, i.e., rejected, response we apply synthetic manipulations through paraphrasing and back-translation via English. The prompts we use to generate the rejected responses are listed in Appendix D. We ensure semantic consistency by using BLEU (Papineni et al., 2002) to filter out pairs of chosen and rejected responses with insufficient similarity. Moreover, to ensure that the linguistic style of the rejected responses differs significantly from the chosen ones, we also filter out pairs with overly high BLEU scores. Based on an empirical analysis of BLEU score distributions, we define the final threshold as $0.15<\textsc{Bleu}(\text{Chosen},\text{Rejected})<0.9$. We also filter out responses shorter than 10 words (often in the style of multiple-choice questions), as they cannot convey enough lexical or syntactic style. The statistics of our two resulting preference datasets are presented in Table 7.

### v2

@@ 1 Introduction
We could make an analogy between these multilingual LLMs and native English speakers who are trying to acquire a new language (Groves et al., 2018; DeVore and Kyle, 2023). Their language notions are built in an English-centric system, and they inevitably bring traces of English habits into other languages when transferring their notions (Papadimitriou et al., 2023; Wendler et al., 2024). Moreover, because of the lack of data in non-English languages, multilingual LLMs are often exposed – either during pre-training or post-training (Yang et al., 2024; Abdin et al., 2024) – to texts translated from English. Both human and machine translated language are known to suffer from translationese artifacts, which set them apart from native content (Bizzoni et al., 2020; Luo et al., 2024). LLMs trained on such data are susceptible to suffer from the same translationese problems.
@@ 2 Related Work
Multilingual analysis of LLMs has recently shown that models trained on unbalanced, English-heavy corpora often rely on English as an internal pivot language. For instance, Wendler et al. (2024) demonstrate that the concept space of LLaMA-2 is more closely aligned with English than with other input languages. More empirically, Papadimitriou et al. (2023) show that multilingual BERT exhibits a bias toward English-like grammatical structures. Despite these findings, there has been no systematic study on how this English-centric tendency affects the linguistic naturalness of multilingual LLM outputs, particularly in open-ended downstream tasks.
@@ 3 Evaluation Metrics for Naturalness
In this section, we introduce a new set of evaluation metrics designed to assess the naturalness of multilingual text generation at the corpus level. While our approach also requires a reference set of natively written texts, it differs from widely used reference-based metrics such as BLEU (Papineni et al., 2002), ROUGE (Lin, 2004), and BERTScore (Zhang* et al., 2020). These sample-level reference metrics often struggle to account for human label variability and uncertainty (Plank, 2022; Giulianelli et al., 2023), particularly in open-ended tasks with multiple valid generations. Additionally, while a single text sample with certain choices of vocabulary or grammatical structure might seem natural, their repeated occurrence across many generations would raise a red flag (Guo et al., 2023). Detecting unnaturalness in single instances can be difficult, but statistical patterns emerge more clearly at the corpus level, serving as stronger indicators (Liang et al., 2024a, b). Our metrics leverage these distributional patterns, offering a broader and more robust perspective on language use across large text collections.
Having highlighted the advantage of a distribution-level perspective, we propose a new definition for language model naturalness. Past studies have defined the naturalness of a single piece of text by asking “could it have been produced by a native speaker?” (Novikova et al., 2016; Groves et al., 2018; Liu et al., 2021). We adapt this definition to the corpus level and define the naturalness of a language model by asking “could the set of texts generated by this language model have been produced by a group of native speakers?”
Our metrics are inspired by divergence measures such as MAUVE (Pillutla et al., 2021; Pimentel et al., 2023), which quantify the information-theoretic divergence between the probability distributions of a language generator and a true natural language distribution. However, our approach differs by not relying on another language model to embed the generated texts. This reduces the risk of introducing intrinsic biases from the chosen embedding model, a crucial consideration in multilingual settings, where such the chosen embedding models are often English-dominated themselves. Our method is also more transparent and interpretable, as it clearly distinguish between two key aspects of linguistic naturalness: syntactic and lexical naturalness. However, our metrics focus exclusively on the linguistic form of the text and do not address semantic aspects. For analyzing semantics, the use of external embedding models may be inevitable.
@@ 3.1 Lexical Naturalness
We propose to measure the lexical naturalness of an LLM by comparing the vocabulary distribution of its generated text with that of human-written text. More specifically, we put forward a lexical naturalness metric based on computing the Jensen-Shannon Divergence (JSD) between the lexical distributions of LLM-generated and human-written text from the same prompts. The JSD provides a symmetric and bounded measure of difference between two distributions without them necessarily sharing the same support, i.e., the tokens in the vocabulary of the LLM in our case. [^2] We process vocabularies at the word level instead of subword/token level to represent more meaningful lexical units. Given the vocabulary distributions $P$ and $Q$ corresponding to human and model outputs, respectively, the JSD is calculated as follows:
@@ 3.2 Syntactic Naturalness
To measure the syntactic naturalness of LLMs generations we leverage the Universal Dependencies (UD) grammar framework (Nivre et al., 2020). UD provides well-defined, theoretically-grounded linguistic structures across different languages, making it suitable for cross-lingual comparisons. Our proposed metric is based on representing each sentence as a dependency tree, where nodes correspond to words and edges specify the dependency relations between them. Additionally, each word is annotated with its corresponding Part-of-Speech (POS) tag as the node label. At a high-level, our metric computes the structural similarity of all pairs of sentences in a corpus, clustering the ones who share similar syntax, allowing us to determine if there is a distributional difference between two groups (LLM-generated text and human-written text, in our case).
@@
Given a set of human-generated sentences $\{s_{i}^{\text{h}}\}_{i=1}^{N_{\text{h}}}$ and model-generated sentences $\{s_{j}^{\text{m}}\}_{j=1}^{N_{\text{m}}}$, we compute the WL kernel similarity between all pairs of dependency trees from these sets. The resulting kernel matrix $\mathbf{K}\in\mathbb{R}^{N_{\text{h}}\times N_{\text{m}}}$ has elements $K_{ij}=K_{\text{WL}}(T_{i}^{\text{h}},T_{j}^{\text{m}})$, where $T_{i}^{\text{h}}$ and $T_{j}^{\text{m}}$ represent the dependency trees corresponding to sentences $s_{i}^{\text{h}}$ and $s_{j}^{\text{m}}$, respectively. This kernel matrix $\mathbf{K}$ captures structural similarities between the dependency trees of human-written and LLM-generated sentences.
@@
To conclude, our proposed syntactic naturalness metric quantifies the syntactic divergence between human and model-generated sentences by examining the distribution of their dependency tree structures. This approach considers both the dependency relationships and the hierarchical arrangement of substructures at multiple levels.
@@ 4 Cross-lingual Analysis of Naturalness
After introducing our new metrics, we proceed with a cross-lingual analysis of the naturalness of multilingual LLMs. This process involves curating a new dataset, selecting the models, and analyzing the benchmark results.
@@ 4.1 Dataset and Evaluation
For our analysis, we require a corpus that is topically aligned across languages while preserving ground-truth naturalness. This means that the texts in each language must be natively written, not translated – whether by humans or models. As a result, the parallel corpora typically used in machine translation research are not suitable. Instead, we construct a new dataset that satisfies our criteria starting from Wikipedia [^3] Wikipedia is available under the CC BY-SA 4.0 license., which offers a wealth of texts that are frequently edited by native speakers and that we can topically-align across languages. Importantly, to minimize cultural bias in our dataset, we select the most-viewed Wikipedia entries across English, Chinese and French. [^4] We use the Wikipedia API to retrieve the top-20 most-viewed entires in each language daily from June 2022 to June 2025. We then take the per-language union of the entries and retain their the cross-language intersection. Our data curation process results in 3,722 Wikipedia entries, each accompanied by descriptions in the three target languages. We discuss the preprocessing of this dataset in Appendix A.
For each entry in the dataset, we instruct the models to complete a straightforward task: generating a description for the given entry in each of the three languages. We select this task to avoid overly constraining the content of the language model outputs, allowing them to generate in a more natural and organic manner. We stress that our focus is on the overall distribution of vocabulary and grammatical structures rather than on individual outputs. The prompts and generation settings used in this task are provided in Appendix B.
@@ 4.3 Results and Analysis
In terms of lexical divergence, English outputs – especially from Mistral-v0.3 – approach human reference values, while larger gaps persist in Chinese and French, with Chinese showing the most significant difference. This indicates that the models are lexically less natural in languages other than English, and the naturalness gap is more pronounced for languages that are typologically distant from English.
@@
Previous benchmark results are generated using temperature sampling with $t=0.6$, which is found to provide the optimal balance between creativity and consistency for multilingual generation tasks (Dubey et al., 2024). As an ablation study, we vary the decoding temperature between 0.3 and 0.9 to examine its impact on the naturalness of model outputs. Our results in Figure 4 demonstrate that for syntactic naturalness, models that are already natural become even better as the decoding temperature increases, whereas models with unnatural outputs show a further increase in unnaturalness. For vocabulary, all models show improved naturalness as the decoding temperature rises. It is also worth noting that the ranking of the models generally remains unchanged across different decoding settings.
@@ 5 Improving Naturalness through Preference Tuning
We now propose an approach to improve the naturalness of language model outputs. Among the various stages of LLM development, preference-based learning is the most effective for refining stylistic features Ivison et al. (2024). Based on this, we focus on aligning models for better naturalness during the preference tuning stage. Among the available preference optimization methods, we opt for DPO due to its simplicity and efficiency.
@@ 5.1 Preference Dataset Construction
We construct the preference datasets starting with SFT (Supervised Fine-Tuning) datasets originally written in the target language. However, we find that non-translated, open-source SFT datasets for non-English languages are almost non-existent. For French, we are unable to locate any such datasets, and for Chinese – despite being the second highest-resource language after English – only a limited number is available. Consequently, we conduct our DPO experiments on Chinese, using two datasets that focus on essay generation and open-domain QA tasks. These tasks are well-suited for stylistic alignment due to their open-ended, creative nature. For essay generation, we use the “composition” split from the Firefly dataset (Yang, 2023), and for open-domain QA, we use the OpenLabel-Chinese Conversations Dataset  [^5] Released under the Apache 2.0 license. (BAAI, 2023).
The original instructions from the SFT datasets are preserved, with the initial response serving as the preferred response. To generate the unnatural, i.e., rejected, response we apply synthetic manipulations through paraphrasing and back-translation via English. The prompts we use to generate the rejected responses are listed in Appendix D. We ensure semantic consistency by using BLEU (Papineni et al., 2002) to filter out pairs of chosen and rejected responses with insufficient similarity. Moreover, to ensure that the linguistic style of the rejected responses differs significantly from the chosen ones, we also filter out pairs with overly high BLEU scores. Based on an empirical analysis of BLEU score distributions, we define the final threshold as $0.15<\textsc{Bleu}(\text{Chosen},\text{Rejected})<0.9$. We also filter out responses shorter than 10 words (often in the style of multiple-choice questions), as they do not convey enough lexical or syntactic style. The statistics of our two resulting preference datasets are presented in Table 7.